Find the general solution of the following differential equations using method of undetermined coefficients: (iv) y''−y =2sinx+2


1
Expert's answer
2022-03-01T18:26:45-0500

Corresponding homogeneous differential equation


"y''\u2212y =0"

Characteristic (auxiliary) equation


"r^2-1=0"

"r_1=1,r_2=-1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^x+c_2e^{-x}"

Find the particular solution of the non homogeneous differential equation


"y_p=A\cos x+B\sin x+C"

"y_p'=-A\sin x+B\cos x"

"y_p''=-A\cos x-B\sin x"

Substitute

"-A\cos x-B\sin x-A\cos x-B\sin x-C=2\sin x+2"

"A=0, B=-1, C=-2"

The general solution of the non homogeneous differential equation is


"y=c_1e^x+c_2e^{-x}-\sin x-2"